OpenStudy (anonymous):

Find y" for y = sin^2(x)

OpenStudy (anonymous):

Did you already find y'?

OpenStudy (anonymous):

so i'm just taking the derivative here right?

OpenStudy (anonymous):

Yes, you are finding the second derivative so just take the derivative of the function and then take the derivative of the derivative of the function :p

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

can u show it please?

OpenStudy (anonymous):

\[y' = 2\sin(x)\cos(x)\]and then\[y'' = \frac{ dy }{ dx } 2\sin(x)\cos(x) = 2(\cos^2(x) - \sin^2(x))\]

OpenStudy (anonymous):

ok